# EdgeOL: Efficient in-situ Online Learning on Edge Devices

### 30 Jan 2024 v1

@@ Abstract.
Emerging applications, such as robot-assisted eldercare and object recognition, generally employ deep learning neural networks (DNNs) models and naturally require: i) handling streaming-in inference requests and ii) adapting to possible deployment scenario changes. Online model fine-tuning is widely adopted to satisfy these needs. However, fine-tuning involves significant energy consumption, making it challenging to deploy on edge devices. In this paper, we propose EdgeOL, an edge online learning framework that optimizes inference accuracy, fine-tuning execution time, and energy efficiency through both inter-tuning and intra-tuning optimizations. Experimental results show that, on average, EdgeOL reduces overall fine-tuning execution time by 82%, energy consumption by 74%, and improves average inference accuracy by 1.70% over the immediate online learning strategy.
11footnotetext: Equal contribution.
@@ 1. Introduction
Deep learning neural networks (DNNs) have gained significant popularity in emerging application domains such as robot-assisted eldercare (do2018rish, ; bemelmans2012socially, ), object recognition (doshi2020continual, ; pellegrini2020latent, ), and wild surveillance (akbari2021applications, ; ke2020smart, ). These applications generally deploy DNN models on energy-constrained edge devices, such as mobile handhelds and internet-of-things (IoT) devices (chen2019deep, ; li2018learning, ; wang2018deep, ; deng2014deep, ). There are two fundamental requirements of deploying DNN models on edge devices: (1) adaptiveness and (2) energy-efficiency. From the perspective of adaptiveness, the DNN applications commonly have streaming-in training data and inference requests over time. This requires model fine-tuning using incoming training data to i) adapt to scenario changes, and meanwhile ii) maximize the average inference accuracy for the streaming-in inference requests (details in Section 2). For example, to maintain high accuracy, an object recognition system needs a timely update of its DNN model when working under different environments or conditions, such as involving new classes of data or instances of existing classes but with new patterns (e.g., strong fog or snow) (belgiovine2022hri, ; antoni2023embedded, ; lomonaco2017core50, ; lomonaco2020rehearsal, ), while keeping the recognition functions online. Regarding energy efficiency, the DNN applications need to optimize energy efficiency since they are often deployed on edge devices with constrained power capacities (wang2019e2, ; yang2017designing, ), such as battery-powered camera drones or mobile phones.
@@
Existing approaches usually employ online learning to ensure model adaptiveness. Immediate online learning is an extreme case of online learning that performs immediate model fine-tuning once new training data arrives (nallaperuma2019online, ; gomes2019machine, ; fontenla2013online, ; hayes2020lifelong, ). As a result, it guarantees high inference accuracy for incoming inference requests since the model is always up-to-date. However, this requires a large amount of computation as well as significant memory overheads from frequent model loading, saving, and system initialization, making it less energy efficient. While fine-tuning (i.e., training) models at a fixed and lower frequency seems to be a reasonable trade-off for accuracy and energy efficiency, it is far from an optimal solution (details discussed in Section 3.2).
@@
In this paper, we aim to achieve both adaptiveness and energy efficiency by proposing a framework called EdgeOL. Our design motivation stems from the observation that there are redundant computation and memory accesses during the fine-tuning stage in immediate online learning. Specifically, we first observe that many fine-tuning rounds contribute little to the inference accuracy. That is, selectively delaying and merging some fine-tuning rounds and reducing the fine-tuning frequency will not hurt the inference accuracy. We call this inter-tuning redundancy. Second, we observe that some layers reach convergence during fine-tuning. Thus, freezing those layers will not affect the inference accuracy. We call this intra-tuning redundancy. Moreover, freezing layers helps avoid over-adaptation to training data and improve the model convergence speed by reducing the number of weights being trained. This allows the streaming-in inference requests to use a robust model with a higher inference accuracy while reducing the fine-tuning time and energy.
@@
    We conduct comprehensive characterization that quantifies the fine-tuning execution time, energy efficiency, and inference accuracy of existing online learning techniques. We reveal that there exist substantial inter-tuning and intra-tuning redundancies that can be optimized to significantly reduce the fine-tuning execution time and energy consumption while improving the inference accuracy.
@@
    We propose EdgeOL that consists of: i) inter-tuning optimization that dynamically and adaptively determines the fine-tuning frequency, and ii) similarity-guided layer freezing for intra-tuning optimization.
@@
    We evaluate EdgeOL using various DNN models and datasets in both computer vision (CV) and natural language processing (NLP) domains. Experimental results show that, compared to immediate online learning in CV domain, EdgeOL saves 82% (76% in NLP domain) of overall fine-tuning execution time and 74% (63% in NLP domain) of energy consumption on average. Furthermore, EdgeOL improves the average inference accuracy of all streaming-in inference requests by 1.70% (1.76% in NLP domain).
@@
    We make a comprehensive evaluation of EdgeOL under different representative scenario change types in online learning, including task change and input distortion. The results show the effectiveness and great generalizability of EdgeOL for different types of application scenarios.
  • $\bullet$
    We demonstrate EdgeOL outperforms state-of-the-art efficient training methods, including layer freezing frameworks i) Egeria and ii) AutoFreeze, iii) sparse training framework RigL, and iv) efficient online learning framework Ekya, even if they have been optimized by our proposed inter-tuning optimization. EdgeOL provides 2.1$\times$, 2.2$\times$, 2.6$\times$, and 2.2$\times$ energy savings, respectively, while delivering 1.76%, 2.10%, 2.17%, and 1.89% higher average inference accuracy.
@@ 2. Background
Online learning. As shown in Figure 1, online learning can be categorized into immediate online learning and fixed-frequency online learning that initiates fine-tuning process after receiving a certain fine-tuning data batches (e.g., two batches in the example in Figure 1). The example involves eight received fine-tuning data batches represented by eight red lines, respectively. The green lines indicate four inference requests. Note that, in practice, inference requests might arrive in bursts (e.g., at $t_{3}$ and $t_{4}$). In immediate online learning, model fine-tuning is triggered once fine-tuning data (i.e., training data) arrives. Thus, the model is fine-tuned eight times in this example. On the other hand, fixed-frequency online learning fine-tunes the model four times in this case. In general, immediate online learning achieves the highest average inference accuracy by frequently updating the model. However, this involves significant overheads from frequent model loading, saving, and system initialization, making it less energy efficient. While fixed-frequency fine-tuning seems to be a reasonable tradeoff between accuracy and energy efficiency, it lacks flexibility and adaptiveness to different cases and is far from an optimal solution (details discussed in Section 3.2).
@@
Average inference accuracy. In online learning, the ongoing model fine-tuning and continuous arrival of inference requests necessitate an evaluation metric to assess the effectiveness of fine-tuning during the entire online learning process. Thus, the average inference accuracy, which is the arithmetic mean of (instantaneous) inference accuracies for all requests, is commonly used to serve the purpose (bhardwaj2022ekya, ). For example, inference requests occur at times $t_{3}$, $t_{4}$, $t_{6}$, and $t_{11}$, with corresponding accuracies $A_{t3}$, $A_{t4}$, $A_{t6}$, and $A_{t_{11}}$. The average accuracy is thus calculated as $(A_{t3}+A_{t4}+A_{t6}+A_{t{11}})/4$.
@@
Scenario change. We categorize the scenario changes into two types: task change and input distortion. For the task change, the initial model is originally trained on a large and general dataset (e.g., ImageNet), then deployed to work on a new dataset (e.g., CIFAR-10) that could contain the data from the classes that are not observed during the previous training. The task change usually happens when a general model needs to be deployed to a more task-specific application (weiss2016survey, ). The input distortion occurs when a model is trained on a particular task/dataset but needs to handle the inference data that are severely distorted, leading to a significant distribution shift (hendrycks2019benchmarking, ). For example, input data for an object recognition system is strongly distorted by natural environments, such as strong fog or snow, which are not considered in the original model training. Conducting online fine-tuning with the new dataset or using the distorted data can effectively mitigate the effect of scenario change and improve accuracy. In our work, we comprehensively evaluate our method for both types of scenario changes.
Figure 3. Overall fine-tuning execution time and energy consumption of nine strategies.
Figure 4. Average inference accuracy using nine strategies.
Figure 3. Overall fine-tuning execution time and energy consumption of nine strategies.
Figure 4. Average inference accuracy using nine strategies.
Figure 5. Accuracy improvement curve of ResNet50 and MobileNetV2 using strategy $S7$.
@@ 3. Challenges and Opportunities
3.1. Experimental Setup
Platform: We use the NVIDIA Jetson Xavier NX and choose the 15W 6-Core power mode with max GPU speed.
Model and dataset: In this section, we employ two popular DNN models, ResNet50 (Res50 for short) (he2016deep, ) and MobileNetV2 (MBV2 for short) (sandler2018mobilenetv2, ). For clarity and simplicity, we use “ImgNet to CF10” as an example test case of scenario change in this section. Specifically, the initial model was well-trained on the ImageNet dataset (deng2009imagenet, ) and will be online fine-tuned and evaluated on CIFAR-10 dataset (krizhevsky2009learning, ) with streaming-in training data (from the training set, containing 50,000 images in total) and inference requests (from the testing set). More experiments are included in Section 5.4 to cover more complex test cases like CORe50 benchmarks (lomonaco2020rehearsal, ), which involve multiple scenario changes. The training batch size is fixed to 8 to avoid out-of-memory errors.
Metrics: We use three metrics for evaluation: overall fine-tuning execution time, overall energy consumption, and average inference accuracy. The overall fine-tuning execution time and energy consumption refer to the total time and energy costs during the entire online learning process. They sum up the fine-tuning execution time and energy consumption of all fine-tuning rounds. The average inference accuracy is the average of accuracies over all inference requests.
Fine-tuning Setting: The arrival granularity of training data is 1 batch each time, and we assume there are a total of 500 inference requests. The arrival rate for both the training data and inference requests follows a Poisson distribution to mimic real application scenarios (mattson2020mlperf, ). We also study different numbers of inference requests and different arrival distributions in Section 5.3. We follow the setting in (lomonaco2017core50, ; lomonaco2020rehearsal, ), where each training sample is used only once in the experiment.
Table 1. The configurations of the fine-tuning strategies.
@@ 3.2. Inter-tuning
We first investigate whether each round of fine-tuning uniformly impacts inference accuracy. As both training data and inference requests are arriving over time, we argue that each fine-tuning round does not contribute equally. As such, those fine-tuning rounds that have little impact on accuracy can be delayed and merged to save time and energy. However, it is non-trivial to determine which rounds are appropriate to be delayed. To this end, we conduct a quantitative study by varying the fine-tuning frequency to reveal these opportunities. Note, the total amount of training data used is not changed as we only delay and merge some fine-tuning rounds. We do not skip using any training data.
@@
We define the fine-tuning frequency as the number of the triggered fine-tuning rounds during a certain period. We only trigger a fine-tuning round after a fixed number of training data batches arrive. With higher fine-tuning frequency, fewer batches are required to trigger a fine-tuning round. As shown in Table 1, we choose nine different fine-tuning frequencies to perform a sweep from immediate online learning (Immed.) to the less frequent fine-tuning strategies $S8$.
@@
3.2.1. Overall fine-tuning execution time and energy consumption
We investigate the overall fine-tuning execution time and energy consumption of the nine strategies during the entire online learning process. All the results are normalized to Immed. As shown in Figure 5, the execution time can be broken down into four parts: 1) system initialization (e.g., model compilation), 2) model loading & saving, 3) data loading, and 4) forward and backward propagation (FP+BP). One can observe that the forward and backward propagation time (grey bar), and data loading time (red bar) are similar across different fine-tuning strategies. This is because the training batch size and the number of total training samples are the same. However, the system initialization time (orange bar), model loading & saving time (blue bar)—the execution overheads that come with each fine-tuning round—grow significantly with increasing fine-tuning frequency. These execution overheads can even dominate the overall execution time and lead to a considerable increase in energy consumption. For example, the energy consumption increases significantly when the fine-tuning frequency increases from $S4$ to Immed. Besides, these overheads vary between models—they account for a higher proportion of the total time in ResNet50 than in MobileNetV2 because ResNet50 (100MB) is 7$\times$ larger than MobileNetV2 (14MB). Additionally, reducing fine-tuning frequency can also effectively save execution time and energy consumption.
3.2.2. Average inference accuracy
While less frequent fine-tuning reduces execution time and energy consumption, it could severely degrade accuracy, which is unacceptable, as shown in Figure 5. Using a higher fine-tuning frequency can help improve the accuracy, but there are diminishing returns as it approaches Immed. The accuracy effectively saturates at a relatively high fine-tuning frequency (e.g., Immed. $\sim S2$ for ResNet50 and Immed. $\sim S1$ for MobileNetV2 in Figure 5). The saturation frequency varies between models due to different convergence speeds and upper-bound accuracies.
3.2.3. Accuracy contribution of each fine-tuning round
Figure 5 displays the model validation accuracy (details of validation accuracy are in the Section 4.1.2) trend over all fine-tuning rounds using strategy $S7$ (other strategies showed similar trends) where fine-tuning is triggered if 500 new training batches arrive (see Table 1). The accuracy improves quickly in early fine-tuning rounds, and the improvement slows in later rounds as the models reach a steady state where the accuracy saturates. This demonstrates the potential to decrease the fine-tuning frequency as the model becomes well-tuned. For instance, some rounds can be delayed and merged with marginal accuracy loss in the steady state. However, it should be noted that the number of fine-tuning rounds where accuracy saturates varies across models, as shown in Figure 5, implying the need for adaptive fine-tuning strategies.

### 16 Mar 2024 v2

@@ Abstract.
Emerging applications, such as robot-assisted eldercare and object recognition, generally employ deep learning neural networks (DNNs) and naturally require: i) handling streaming-in inference requests and ii) adapting to possible deployment scenario changes. Online model fine-tuning is widely adopted to satisfy these needs. However, an inappropriate fine-tuning scheme could involve significant energy consumption, making it challenging to deploy on edge devices. In this paper, we propose EdgeOL, an edge online learning framework that optimizes inference accuracy, fine-tuning execution time, and energy efficiency through both inter-tuning and intra-tuning optimizations. Experimental results show that, on average, EdgeOL reduces overall fine-tuning execution time by 64%, energy consumption by 52%, and improves average inference accuracy by 1.75% over the immediate online learning strategy.
copyright: none11footnotetext: Equal contribution.
@@ 1. Introduction
With the exceptional performance, Deep learning neural networks (DNNs) have gained significant popularity in emerging application domains such as robot-assisted eldercare (do2018rish, ; bemelmans2012socially, ), object recognition (doshi2020continual, ; pellegrini2020latent, ), and wild surveillance (akbari2021applications, ; ke2020smart, ). These cutting-edge applications generally deploy DNN models on energy-constrained edge devices, such as robots and internet-of-things (IoT) devices (lesort2020continual, ; zhao2021general, ; chen2019deep, ; li2018learning, ; wang2018deep, ).
@@
There are two fundamental requirements of deploying DNN models on edge devices: (1) adaptiveness and (2) energy-efficiency. From the perspective of adaptiveness, the DNN applications commonly have streaming-in training data and inference requests over time. This requires model fine-tuning using incoming training data to i) adapt to scenario changes, and meanwhile ii) maximize the average inference accuracy for the streaming-in inference requests (details in Section 2). For example, to maintain high accuracy, an object recognition system needs a timely update of its DNN model when working under different environments or conditions, such as involving new classes of data or instances of existing classes but with new patterns (e.g., different illumination conditions, background, and occlusion) (lomonaco2017core50, ; lomonaco2020rehearsal, ; she2020openloris, ; roy2023cl3, ; hao2023enhanced, ) while keeping the recognition functions online. Regarding energy efficiency, the DNN applications need to optimize energy efficiency since they are often deployed on edge devices with constrained power capacities (wang2019e2, ; yang2017designing, ; wu2021enabling, ), such as battery-powered robots, mobile phones, or IoT devices.
@@
Existing approaches usually employ online learning to ensure model adaptiveness. Immediate online learning is an extreme case of online learning that performs immediate model fine-tuning once new training data arrives (nallaperuma2019online, ; gomes2019machine, ; fontenla2013online, ; hayes2020lifelong, ). As a result, it guarantees high inference accuracy for incoming inference requests since the model is always up-to-date. However, this requires a large amount of computation as well as significant overheads from frequent model loading, saving, and system initialization like model compilation, making it less energy efficient. On the other hand, fine-tuning (i.e., training) models at a fixed and lower frequency seems to be a reasonable trade-off for accuracy and energy efficiency. But it is still far from an optimal solution (details are discussed in Section 3.1).
@@
In this paper, we propose EdgeOL, an online learning framework for edge devices, aiming to achieve both adaptiveness and energy efficiency. Our design motivation stems from the observation that there are redundant computation and memory accesses during the model fine-tuning stage in existing online learning approaches. Specifically, we first observe that many fine-tuning rounds contribute little to the inference accuracy. That is, selectively delaying and merging some fine-tuning rounds and reducing the fine-tuning frequency will not hurt the inference accuracy. We call this inter-tuning redundancy. Second, we observe that some layers gradually reach convergence during fine-tuning. In this case, freezing those converged layers will not affect the inference accuracy. We call this intra-tuning redundancy. Moreover, freezing layers helps to avoid over-adaptation to the training data and improve the model convergence speed by reducing the number of weights being trained. This allows the streaming-in inference requests to use a robust model with higher accuracy while reducing the fine-tuning time and energy consumption.
@@
    We conduct a comprehensive characterization that quantifies the fine-tuning execution time, energy consumption, and inference accuracy of existing online learning approa- ches. We reveal that there exist substantial inter-tuning and intra-tuning redundancies that can be optimized to significantly reduce the fine-tuning execution time and energy consumption while improving the inference accuracy.
@@
    We propose EdgeOL framework that consists of: i) inter-tuning optimization that dynamically and adaptively determines the fine-tuning frequency, and ii) similarity-guided layer freezing for intra-tuning optimization.
@@
    We evaluate EdgeOL using various DNN models and datasets in both computer vision (CV) and natural language processing (NLP) domains. Experimental results show that, compared to immediate online learning in CV domain, EdgeOL saves 64% (67% in NLP domain) of overall fine-tuning execution time and 52% (54% in NLP domain) of energy consumption on average. Furthermore, EdgeOL improves the average inference accuracy of all streaming-in inference requests by 1.75% (1.52% in NLP domain).
@@
    We demonstrate EdgeOL outperforms state-of-the-art efficient training methods, including layer freezing frameworks i) Egeria and ii) SlimFit, iii) sparse training framework RigL, and iv) efficient online learning framework Ekya, even if they have been optimized by our proposed inter-tuning optimization. EdgeOL provides 2.1$\times$, 2.2$\times$, 2.8$\times$, and 2.0$\times$ energy savings, respectively, while delivering 1.78%, 2.18%, 2.33%, and 1.50% higher accuracy.
@@ 2. Background
Scenario change. The deployment scenario of an already-in-use model may change over time as the user usage scenario evolves (logacjov2021learning, ; liu2021lifelong, ; irfan2021lifelong, ; pinto2016curious, ; ma2023cost, ). These changes can generally be classified into two types, the introduction of i) instances of existing data classes but with new feature patterns (hendrycks2019benchmarking, ; she2020openloris, ; lomonaco2020rehearsal, ) and ii) new classes of data (lomonaco2017core50, ; roy2023cl3, ; hao2023enhanced, ). Instances with new feature patterns refer to scenarios where the model encounters variations (e.g., different illumination conditions, background, and occlusion) in previously recognized data classes. These variations could be due to changes in environmental conditions, user behavior, or other factors that alter the appearance or characteristics of the data. On the other hand, the introduction of new classes of data presents a different challenge, where the model must learn to identify classes that were completely absent previously. In our work, we comprehensively evaluate our method for both types of scenario changes.
Online learning. Conducting online learning can effectively mitigate the effect of scenario change and improve accuracy, which is essential for models to function effectively in the ever-changing real-world environment(shmelkov2017incremental, ; bhardwaj2022ekya, ; rebuffi2017icarl, ). In online learning, the fine-tuning data is not well-prepared all at once but rather constantly streaming in, with new data arriving continuously throughout each scenario (wu2021enabling, ; pinto2016curious, ; lomonaco2017core50, ). Figure 1 illustrates the immediate online learning and an example of fixed-frequency online learning that initiates fine-tuning process after receiving a certain fine-tuning data batches (e.g., two batches in the example in the figure). We assume two scenario changes in the figure, indicated by the red dotted arrows right before $t_{1}$ and $t_{12}$. It involves eight received fine-tuning data batches represented by eight red lines, respectively. The green lines indicate four inference requests. Note that, in practice, inference requests might arrive in bursts (e.g., at $t_{7}$ and $t_{8}$). In immediate online learning, model fine-tuning is triggered once fine-tuning data (i.e., training data) arrives. Thus, the model is fine-tuned eight times in this example. On the other hand, fixed-frequency online learning fine-tunes the model four times in this case. In general, immediate online learning achieves the highest average inference accuracy by frequently updating the model. However, this involves significant overheads from frequent model loading, saving, and system initialization (e.g., model compilation), making it less energy efficient. While fixed-frequency fine-tuning seems to be a reasonable tradeoff between accuracy and energy efficiency, it lacks flexibility and adaptiveness to different cases and is far from an optimal solution (details discussed in Section 3.1).
@@
Average inference accuracy. In online learning, the ongoing model fine-tuning and continuous arrival of inference requests necessitate an evaluation metric to assess the effectiveness of fine-tuning during the entire online learning process. Thus, the average inference accuracy, which is the arithmetic mean of (instantaneous) inference accuracies for all requests, is commonly used to serve the purpose (bhardwaj2022ekya, ). For example, inference requests occur at times $t_{2}$, $t_{7}$, $t_{8}$, and $t_{11}$, with corresponding accuracies $A_{t2}$, $A_{t7}$, $A_{t8}$, and $A_{t_{11}}$. The average accuracy is thus calculated as $(A_{t2}+A_{t7}+A_{t8}+A_{t{11}})/4$.
@@ 3. Challenges and Opportunities
In this section, we employ two popular DNN models Res- Net50 (he2016deep, ) and MobileNetV2 (sandler2018mobilenetv2, ), and use the NC (New Class) benchmark in the widely-used CORe50 dataset (lomonaco2017core50, ) as an example for testing. There are 9 scenarios in total in this benchmark and the scenarios appear one after one, each of which introduces new classes of data on top of the existing classes. The model is originally well-trained using the training data in the first scenario and will be online fine-tuned with corresponding training data and serve inference requests in each subsequent scenario (i.e., scenario 2$\sim$9). Both the training data and inference requests arrive continuously over time. Please refer to Section 5.1 for details of the experimental setup.
Table 1. The configurations of different online learning strategies.
@@ 3.1. Inter-tuning
Recall that the online learning approach tunes a model whenever sufficient training data arrives (e.g., a batch of training data). However, this timely and immediate fine-tuning approach consumes time and energy. We conduct a quantitative study by varying the fine-tuning frequency.
@@
We consider the fine-tuning frequency as the number of the triggered fine-tuning rounds during a certain period, where each round is triggered after a fixed number of training data batches arrive. Therefore, with higher fine-tuning frequency, fewer batches are required to trigger a fine-tuning round. As shown in Table 1, we choose seven different fine-tuning frequencies, from immediate online learning (Immed.) to the less frequent fine-tuning strategies $S6$. Note that, in this experiment, the total amount of used training data is not changed as we only delay and merge some fine-tuning rounds, where a round will involve more training data. We do not skip using any training data.
@@
Figure 3. Overall fine-tuning execution time and energy consumption of seven strategies.
Figure 4. Average inference accuracy using seven strategies.
Execution time, energy, and accuracy. We investigate the overall fine-tuning execution time and energy consumption of the seven strategies during the entire online learning process. The results are normalized to Immed. As one can observe from Figure 3, the fewer number of fine-tuning rounds (i.e., from Immed to S6), the more savings in overall execution time and energy consumption. This is because frequently triggering fine-tuning introduces significant overheads such as model loading, saving, and system initialization (e.g., model compilation). However, the savings in time and energy come with accuracy degradation, as shown in Figure 4. Therefore, an important question we seek to answer is that How can we reduce the fine-tuning frequency while maintaining the accuracy?
Accuracy improvement by each fine-tuning round. To answer the above question, Figure 5 takes strategy $S5$ as an example (other strategies showed similar trends) and shows the model validation accuracy [^1] Details of validation accuracy are defined in the Section 4.1.2 over fine-tuning rounds in the two consecutive scenarios. One can make the following observations from the results. First, as expected, there is a significant accuracy drop when there is a scenario change. Second, after the scenario change, the validation accuracy improves quickly in early fine-tuning rounds and saturates in later rounds. This demonstrates that not every fine-tuning round contributes significantly to the accuracy. Therefore, one can potentially delay and merge some of the fine-tuning rounds without affecting the accuracy. Third, the number of fine-tuning rounds where accuracy saturates varies across models, as indicated by the two models in Figure 5, implying the need for an adaptive approach to determine the fine-tuning rounds that could be delayed and merged.
Figure 5. Accuracy improvement curve of ResNet50 and MobileNetV2 in two consecutive scenarios using strategy $S5$.